OpenStudy (anonymous):

Charlie wants to buy a $800 TV in 8 months. How much should he invest now at 10% simple interest to have the money in 8 months?

OpenStudy (anonymous):

I = PRT, the simple interest formula. R = 10%, P is what we must find, T is 8/10 of a year, I = interest on P earned in 8 months. But, P + I = $800 So, P + I = P + PRT = P(1 + RT) So, P = 800/(1+RT) = 800/(1 +0.1x0.8) = $740.74 Check! What will $740.74 grow to at 10% interest for 8 months? I = PRT = 740.74 x 0.1 x 0.8 = 59.26 So &40.74 plus interest of 59.26 grows to 740.74 + 59.26 = $800